OpenStudy (chihiroasleaf):

Given that two circles touch y-axis and line y=(1/sqrt(3)) x. if the centre of both circles lie on the line y = sqrt(3) x what is the distance between the centre of both circles? any hint? :)

OpenStudy (chihiroasleaf):

sorry..., I mistype the question..., the centre of both circles lie on the line y = sqrt (3)

OpenStudy (sirm3d):

hmm, i got 4 for an answer.

OpenStudy (sirm3d):



hartnn (hartnn):

two circles with radius 1 and 3, so distance between centres = sum of radius =4,( as both the circles touch y-axis) could u find radius of each circle ?

hartnn (hartnn):

the line has slope of 1/sqrt 3, so the angle that the line makes with x-axis =30 degrees

hartnn (hartnn):

so u get 2 , 30-60-90 triangles, from where u can get radius of both circles

OpenStudy (sirm3d):

the coordinate of the first center is (-b,sqrt(3)), then use the distance of a point from a line (y=sqrt(3)x) which is also b.

OpenStudy (sirm3d):

oops, i think i made some mistake in my algebra.

OpenStudy (sirm3d):

my algbra is correct. the answer is still 4. what did you get, @hartnn ?

hartnn (hartnn):

even i get 4.

hartnn (hartnn):